OpenStudy (anonymous):

(5a^4/2b^8)2 simplifyplease

OpenStudy (anonymous):

\[\Large (\frac{5a^4}{2b^8})^2\] \[\Large \implies \frac{(5a^4)^2}{(2b^8)^2}\] \[\Large \frac{5^2(a^4)^2}{2^2(b^8)^2}\] can you do this now??

OpenStudy (anonymous):

@cdelomas

OpenStudy (anonymous):

ok is the lsoe e answer

OpenStudy (anonymous):

u need me to simplify it more ??

OpenStudy (anonymous):

The question just saidto simplify the problem

OpenStudy (anonymous):

\[\Large \implies \frac{25a^8}{4b^{16}}\]

OpenStudy (anonymous):

ok